The length of a rectangle is three times it's width.
If the perimeter of the rectangle is 72 yards, find it's area.
Need help please!

Answer by checkley79(3341)

L=3W
2L+2W=72
2(3W)+2W=72
6W+2W=72
8W=72
W=72/8
W=9 YARDS FOR THE WIDTH
L=3*9=27 ANS. FOR THE LENGTH
AREA=LW
AREA=27*9
AREA=243 SQUARE YARDS IS THE ANSWER.